The ABC bank has a deposit base on which it holds excess reserves of $42 Million at a reserve rate which relates to a money multiplier (MM) of 5. On January 2, 2010 Mwen-Mwen, a customer of ABC bank issued a $20 Million check drawn on its account, to Bushrod Manufacturing Company (BMC), which it deposited into its account with Satiah Bank.
1
Expert's answer
2021-09-24T11:27:10-0400

Question: find the reserve ratio

Money multiplier"=\frac{1}{Reserve\space ratio}\\5=\frac{1}{Reserve\space ratio}\\=Reserve\space ratio=\frac{1}{5}=20\%"